Suppose the revenue and cost function of a hypothetical firm producing y commodity is given by:

               R(y) =45y-1/2y2

               C(y) = y3-8y2+57y+2

a)        Find the level of output that maximizes revenue of the firm.

b)        At what level of output the average variable cost of the firm is minimized?

c)        Calculate the output level that maximizes profit and maximum profit of the firm.


1
Expert's answer
2021-04-21T15:02:17-0400
"MR=45-y"

"MC=3y^2-16y+57"

"y_{max}=1"

"y_{min}=4"

"\pi= R(y)-C(y)=(45-0.5)-(1-8+57+2)=-7.5"